OpenStudy (anonymous):

A plane is flying east at 115 m/s. the wind accelerates it at 2.88m/s^2 directly northwest. after 25.0 s what is the velocity (magnitude and direction) of the plane?

OpenStudy (shamim):

i wanna try it

OpenStudy (anonymous):

Ok

OpenStudy (shamim):

accelereation has 2 perpendicular component

OpenStudy (shamim):

1 component along west is\[a _{w}=2.88 \cos 45\]

OpenStudy (shamim):

another component along north is\[2.88 \sin 45\]

OpenStudy (shamim):

now after 25 sec the velocity along east wil\[v _{e}=115-2.88 \cos 45 \times 25\]l b

OpenStudy (shamim):

after 25 sec the velocity along north will b\[v _{n}=2.88 \sin 45 \times 25\]

OpenStudy (shamim):

so the magnitude of velocity after 25 sec \[v=\sqrt{v _{e}^{2}+v _{n}^{2}}\]

OpenStudy (shamim):

and the angle of velocity with north will b\[\tan \theta =\frac{ v _{e} }{ v _{n} }\]

OpenStudy (shamim):

m i right?????????

OpenStudy (anonymous):

Yeah